# Testing Monotonicity of Real-Valued Functions on DAGs

### v1

@@ Abstract
We study monotonicity testing of real-valued functions on directed acyclic graphs (DAGs) with $n$ vertices. For every constant $\delta>0$, we prove a $\Omega(n^{1/2-\delta}/\sqrt{\varepsilon})$ lower bound against non-adaptive two-sided testers on DAGs, nearly matching the classical $O(\sqrt{n/\varepsilon})$-query upper bound. For constant $\varepsilon$, we also prove an $\Omega(\sqrt{n})$ lower bound for randomized adaptive one-sided testers on explicit bipartite DAGs, whereas previously only an $\Omega(\log n)$ lower bound was known. A key technical ingredient in both lower bounds is positive-matching Ruzsa–Szemerédi families. On the algorithmic side, we give simple non-adaptive one-sided testers with query complexity $O(\sqrt{m\,\ell}/(\varepsilon n))$ and $O(m^{1/3}/\varepsilon^{2/3})$, where $m$ is the number of edges in the transitive reduction and $\ell$ is the number of edges in the transitive closure. For constant $\varepsilon>0$, these improve over the previous $O(\sqrt{n/\varepsilon})$ bound when $m\ell=o(n^{3})$ and $m=o(n^{3/2})$, respectively.
@@ 1 Introduction
Testing monotonicity on structured domains is a central problem in property testing and a canonical example in the sublinear-time model; see, e.g., [17, 23, 8, 22, 29]. It was first studied on the Boolean hypercube and product domains and has since been extended to richer product structures and general posets/graph families [2, 7, 11, 16, 19, 21, 25, 28]. A natural way to model posets is via reachability in directed acyclic graphs (DAGs). In this work we study the massively parameterized model [29, 23], where the entire DAG is given explicitly.
@@ 1.1 Problem and background
Let $G=(V,E)$ be a DAG with $n=|V|$ vertices and transitive closure $\mathrm{TC}(G)$. A function $f:V\to\mathbb{R}$ is monotone if $f(u)\leq f(v)$ whenever $u\leadsto v$, i.e., for every edge in $\mathrm{TC}(G)$. Given full access to $G$, a proximity parameter $\varepsilon>0$, and oracle access to $f$, the goal is to distinguish monotone functions from those that are $\varepsilon$-far from monotonicity with constant success probability. We do not charge for preprocessing or computations on $G$; only queries to $f$ are counted. We consider both non-adaptive and adaptive testers, with one-sided or two-sided error; see Section 2 for definitions. A tester is non-adaptive if all queries are fixed in advance (adaptive otherwise), and one-sided if it always accepts monotone functions (two-sided otherwise). Much of the monotonicity-testing literature on both upper and lower bounds focuses on product domains such as the hypercube and hypergrid [5, 10, 11, 12, 13, 14, 15, 16].
@@
queries, where $\ell=|E(\mathrm{TC}(G))|$ is the number of comparable pairs. They also show non-adaptive two-sided lower bounds, for Boolean ranges, based on Ruzsa–Szemerédi-type constructions [19], yielding a bound of $n^{\Omega(1/\log\log n)}$. This is still far from the abovementioned upper bound, leaving a large gap. Beyond the Fischer et al. result on general posets, most sharper lower bounds are proved for product domains (in particular, hypercubes/hypergrids), and many are formulated in the Boolean-range setting.
@@ 1.2 Our results
Our main contribution is a sharper picture of query complexity for monotonicity testing on DAGs: we strengthen lower bounds in both non-adaptive and adaptive models, and complement them with improved non-adaptive one-sided upper bounds in sparse regimes. The quantitative statements and their comparison with prior work are summarized in Table 1.
@@
In the adaptive regime, logarithmic lower bounds are known already on the total order (equivalently, path graph) [17]. For constant $\varepsilon$ and infinitely many $n$, we prove an $\Omega(\sqrt{n})$ lower bound on explicit bipartite $n$-vertex DAGs for adaptive one-sided testers (Theorem 4.28).
@@
  • A hybrid sampling bound: there is a tester with $q=O\!\left(\sqrt{m\,\ell}/(\varepsilon n)\right)$ queries (Theorem 5.1).
@@
These upper bounds can improve the $O(\min\{\sqrt{n/\varepsilon},\ell/(\varepsilon n)\})$ bound in sparse or structured regimes, complementing the lower bounds above. They explicitly exploit both the transitive reduction and the transitive closure, and are conceptually related to reachability sparsification and TC-spanner-based testers and reconstruction on posets [6, 7, 28, 31].
@@ 1.3.1 Non-adaptive two-sided lower bound
The classical lower-bound approach for monotonicity testing on general posets, pioneered by Fischer et al. [18], is built around Ruzsa–Szemerédi (RS) graphs: one packs many large induced matchings and uses them to generate hard YES/NO instances. The relevance of matchings is that violated pairs in a matching are vertex-disjoint, so repairing all of them requires changing at least one endpoint per pair, which directly yields a distance-from-monotonicity lower bound. For query lower bounds, matchings also localize evidence: a tester typically needs to query both endpoints of one matched pair to observe informative violations. Inducedness then plays a dual role there. It is what makes the violation pattern “clean” for Boolean-valued functions (extra comparabilities among matching endpoints would immediately create additional violations), and at the same time it becomes the bottleneck: extremal graph theory severely limits how many linear-size induced matchings can coexist in a single graph, which is why RS-based templates stall far below the $\tilde{\Theta}(\sqrt{n})$ regime.
Our first conceptual step is to identify the right analogue of induced matchings for real-valued functions. On bipartite DAGs (all edges oriented left-to-right), the question “can a matching be realized as exactly the set of violated edges?” has a precise answer: this holds if and only if the matching is positive (Section 3.1). Informally, “positive” means there is a vertex weight assignment $w$ such that $w(\ell)+w(r)>0$ exactly on matching edges and $w(\ell)+w(r)\leq 0$ on all other edges. Equivalently, positivity excludes alternating cycles/closed walks on the matched vertices (Theorem 2.3). This motivates our central combinatorial abstraction: positive-matching RS (PMRS) families, obtained by relaxing RS from induced matchings to positive matchings (Section 3.2). This relaxation is tailored to the real-valued setting and bypasses the inducedness barrier that constrains the Fischer et al. framework.
@@
Once a PMRS decomposition is available, the lower bound becomes essentially information-theoretic. Given $s$ edge-disjoint positive matchings $M_{1},\dots,M_{s}$ (each of linear size), we hide a random index $i\in[s]$ and construct two distributions over functions. The YES distribution is monotone, built from a base monotone function with a constant margin on non-matching edges (so small random noise does not create new violations), while the NO distribution is $\varepsilon$-far because it violates a constant fraction of edges in the hidden matching $M_{i}$ (Section 3.3). Crucially, the two distributions have identical single-vertex marginals; they differ only in a local correlation between the two endpoints of a matching edge. Therefore, a non-adaptive algorithm can obtain distinguishing information only if it queries both endpoints of some edge in $M_{i}$. Since the matchings are edge-disjoint, a $q$-query set can fully cover at most $O(q^{2})$ edges, and the probability of hitting the hidden matching index is at most $O(q^{2}/s)$, yielding the $\Omega(\sqrt{s})$ lower bound (Theorem 3.8). A useful technical fact (as with induced matchings) is that positivity is closed under taking submatchings (Section 3.1), which lets us refine each large matching into many smaller positive matchings and cleanly extract the $1/\sqrt{\varepsilon}$ dependence (Sections 3.5 and 3.5).
@@
The remaining challenge is to build explicit graphs with many large edge-disjoint positive matchings. This is precisely where the departure from induced matchings matters: unlike RS graphs, PMRS packings exhibit different extremal behavior and admit simple polynomial-size constructions. Our construction (Section 3.4) uses vertex labels $(x,z)$ with $x\in[N]^{k}$ and an auxiliary “height” coordinate $z\in[N^{2}]$. Each matching is indexed by a shift vector $a$ and consists of translated edges $(x,z)\mapsto(x+a,\,z+\|a\|_{2}^{2})$. The nontrivial point is positivity: we design a weight function $w_{a}$ so that for an edge belonging to shift $b$, the positivity-witness sum $w_{a}(\ell)+w_{a}(r)$ (for edge endpoints $\ell,r$) collapses to a quantity of the form $1/2-\|b-a\|_{2}^{2}$, which is positive exactly when $b=a$ (Theorem 3.9). This “quadratic lift” is the key mechanism that separates different matchings by a single linear functional and enables $s=\Omega(n^{k/(k+2)})$ edge-disjoint linear-size positive matchings. Combined with the PMRS $\Rightarrow$ lower-bound template above, this yields our near-$\sqrt{n/\varepsilon}$ non-adaptive two-sided lower bound.
@@ 1.3.2 Adaptive one-sided lower bound
Softly planted violations and the witness principle.
Moving from non-adaptive to adaptive testers, and from two-sided to one-sided error, changes the lower-bound game qualitatively. In the bipartite-DAG setting (all edges oriented left-to-right), one-sidedness comes with a very strong “witness principle” (Section 4): if the tester has not queried a violated edge, then the partial transcript can always be extended to a globally monotone function. In other words, a one-sided tester is not “estimating distance” indirectly; it is literally searching for a witnessed violation. The challenge is that adaptivity seems to help precisely with such search problems: after observing a few values, the tester can steer future queries toward edges that are more likely to be violated.
Section 4’s strategy is to turn this apparent strength of adaptivity into a liability by embedding the search problem into an indexing problem. As in the PMRS-based non-adaptive lower bound, we start from a scaffold graph containing many edge-disjoint positive matchings $M_{1},\dots,M_{s}$, and we hide a random index $I\in[s]$. The goal is to generate a hard distribution over functions such that (i) with high probability the sampled function is $\varepsilon$-far from monotone because it violates almost all edges of the hidden matching $M_{I}$, (ii) edges outside $M_{I}$ are “almost never” violated (so the tester cannot reject without learning something about $I$), and (iii) even an adaptive algorithm learns only a tiny amount of information about $I$ per query. The last two requirements are exactly where naive “plant a matching of violations” constructions break: once we condition on an adaptive prefix, the remaining unqueried variables are no longer independent, and a tester might try to create (or expose) violations on non-matching edges by strategically revealing endpoints.
The key conceptual move is to replace a combinatorial YES/NO gadget by a soft, analytically controlled prior: a truncated, strongly log-concave Gibbs distribution $\{\mu_{i}\}_{i\in[s]}$ over real-valued labelings $f$. The Hamiltonian has three roles. A quadratic “pinning + smoothing” backbone
@@
enforces global strong convexity, which later yields stability of conditional marginals under arbitrary transcript conditioning. On top of this, hinge-square potentials implement a “margin-separated” monotonicity pattern: for edges $e\notin M_{i}$ we add a penalty $\beta(f(\ell)-f(r)+\gamma)_{+}^{2}$, which energetically prefers $f(\ell)-f(r)\leq-\gamma$ (monotone with slack), while for edges $e\in M_{i}$ we add $\beta(\gamma-(f(\ell)-f(r)))_{+}^{2}$, which prefers $f(\ell)-f(r)\geq\gamma$ (a robust violation). Finally, we truncate to a box $B,B]^{V}$ to keep all Lipschitz/concentration arguments uniform. The result is a “softly planted” violated matching: violations are not imposed as hard constraints, but they are so energetically favored that they occur with overwhelming probability, and crucially this preference survives conditioning on adaptively revealed coordinates.
Two technical bottlenecks then dominate the proof, and Section 4 resolves them via two lines of analysis: spurious control and information leakage, explained below.
@@
Even if a fresh non-matching edge is very unlikely to be violated unconditionally, an adaptive tester might try to bias it by first querying one endpoint and then deciding whether to query the other. The analysis therefore must be prefix-conditioned: after any transcript prefix, when the tester “closes” a new edge, the conditional probability of witnessing a violation on an edge outside $M_{I}$ must still be tiny. This is handled by a careful one-dimensional reduction on the difference coordinate $d=f(\ell)-f(r)$ for an edge $e=(\ell,r)$: after conditioning on everything except the two endpoints, the induced density of $d$ is log-concave with explicit strong curvature coming from the quadratic backbone. The hinge term sets a target point ($\gamma/2$ for matching edges and $\gamma/2$ for non-matching edges), and strong convexity translates “the derivative at the target has the right sign” into an exponentially small tail probability of landing on the wrong side (violated vs. non-violated).
The subtlety is the external drift contributed by all other incident constraints: under adaptive conditioning, this drift is random and depends on the revealed neighborhood. Section 4 controls this drift at the two critical points $d=\pm\gamma/2$ with a high-probability argument: for $C_{4}$-free scaffolds, the relevant neighborhood variables become independent after conditioning on the endpoints, and strong log-concavity gives subgaussian tails for the drift (the “drift-tail” bound in Section 4.2). This is paired with a separate “mean-control” step to ensure the conditional mean of the drift is also small; in the final instantiation on our explicit dense scaffold (Section 4.7), this is the most delicate analytic point.
There is a second, purely one-sided obstacle: the box truncation introduces boundary effects, and an adaptive tester might try to push queried values near $\pm B$, where the feasible interval of $d$ shrinks and the clean tail bound could fail. The fix is an “inward drift” phenomenon (Sections 4.1 and 4.1): when $\beta>\lambda$, the hinge potentials create a deterministic force that pushes single-site conditionals away from the dangerous boundary strips. This yields a strategy-independent bound on the probability that any query ever lands near $\pm B$ (Section 4.3), and hence bounds the probability that an adaptive strategy can exploit truncation to manufacture spurious witnesses. Putting these together yields an adaptive spurious-control theorem (Section 4.3): with high probability, any $q$-query adaptive strategy witnesses no violated edge outside the hidden matching unless it hits an exponentially unlikely boundary event.
@@
Even if spurious violations are suppressed, an adaptive tester could hope to learn the hidden index $I$ from the queried values, and then target an edge of $M_{I}$. Section 4 controls this via a mutual-information upper bound $I(I;T)$, where $T$ is the full adaptive transcript. The core technical ingredient is a “Green-function comparison” inequality for uniformly convex Gibbs measures (Section 4.4): for any prefix conditioning $\tau$, the KL divergence between the one-site conditional distributions under $\mu_{i}(\cdot\mid\tau)$ and $\mu_{j}(\cdot\mid\tau)$ is bounded by a weighted sum of local discrepancies, with weights given by entries of $(\alpha I+2\lambda L_{\tau})^{-1}$, where $L_{\tau}$ is the Laplacian on the still-unqueried vertices after conditioning on the prefix $\tau$. Unlike Dobrushin-type arguments, this comparison needs no global mixing/uniqueness regime; it leverages only the strong convexity provided by the quadratic backbone and survives arbitrary adaptive conditioning. A simple row-sum bound on this Green function then gives a clean per-query leakage scale $\kappa_{Q}$, leading to the budgeted bound $I(I;T)\leq q\kappa_{Q}$ uniformly over all adaptive strategies with $q\leq Q$ (Section 4.4).
With spurious-control and leakage in place, the lower bound becomes information-theoretic in a way that is tailor-made for one-sided testing. Because the matchings are edge-disjoint, any witnessed violated edge in $\bigcup_{i}M_{i}$ uniquely identifies the hidden index $I$. By the one-sided witness principle, rejection implies the transcript contains a witnessed violated edge; by spurious-control, with high probability this violated edge lies in $M_{I}$. Thus any successful tester yields a decoder that recovers $I$ from $T$ with constant probability, and a standard Fano argument (Section 4.5) forces $I(I;T)\gtrsim\log s$. Comparing $\log s$ to the leakage upper bound $q\kappa_{Q}$ produces a fixed-budget barrier (Theorem 4.22): below a certain query budget, no adaptive one-sided tester can simultaneously (a) keep spurious witnesses rare and (b) accumulate enough information to find a true matching-edge violation reliably.
@@
Section 4.7 instantiates this framework on an explicit dense scaffold: a $C_{4}$-free, almost-regular subgraph of the shift-matching construction (Section 4.7) with $s=\Theta(\sqrt{n})$ disjoint positive matchings. In this regime the auxiliary probabilities (matching-edge non-violation, spurious violations, boundary failures, and mean-drift failures) can all be driven superpolynomially small by setting the margin $\gamma$ to grow like $\sqrt{\log Q}$, while the per-query leakage remains on the order of $(\log Q)/\sqrt{n}$. Choosing $Q\asymp\sqrt{n}$ then makes the information barrier bite and yields an $\Omega(\sqrt{n})$ adaptive one-sided lower bound for a constant $\varepsilon$ (Theorem 4.28).

### v2

@@ Abstract
We study monotonicity testing of real-valued functions on directed acyclic graphs (DAGs) with $n$ vertices. For every constant $\delta>0$, we prove a $\Omega(n^{1/2-\delta}/\sqrt{\varepsilon})$ lower bound against non-adaptive two-sided testers, nearly matching the classical $O(\sqrt{n/\varepsilon})$-query upper bound. For constant $\varepsilon$, we also prove an $\Omega(\sqrt{n})$ lower bound for adaptive one-sided testers, whereas previously only an $\Omega(\log n)$ lower bound was known. Both lower bounds rely on positive-matching Ruzsa–Szemerédi families. On the algorithmic side, we give simple non-adaptive one-sided testers with query complexity $O(\sqrt{m\,\ell}/(\varepsilon n))$ and $O(m^{1/3}/\varepsilon^{2/3})$, where $m$ is the number of edges in the transitive reduction and $\ell$ is the number of edges in the transitive closure. For constant $\varepsilon>0$, these improve over the previous $O(\sqrt{n/\varepsilon})$ bound when $m\ell=o(n^{3})$ and $m=o(n^{3/2})$, respectively.
@@ 1 Introduction
Testing monotonicity on structured domains is a central problem in property testing and a standard example in the sublinear-time model; see, e.g., [17, 23, 8, 22, 29]. It was first studied on the Boolean hypercube and product domains and has since been extended to richer product structures and general posets/graph families [2, 7, 11, 16, 19, 21, 25, 28]. One way to model posets is via reachability in directed acyclic graphs (DAGs). In this work we study the massively parameterized model [29, 23], where the entire DAG is given explicitly.
@@ 1.1 Problem and background
Let $G=(V,E)$ be a DAG with $n=|V|$ vertices. A function $f:V\to\mathbb{R}$ is monotone if $f(u)\leq f(v)$ whenever $(u,v)\in E$. Given full access to $G$, a proximity parameter $\varepsilon>0$, and oracle access to $f$, the goal is to distinguish monotone functions from those that are $\varepsilon$-far from monotonicity with constant success probability. We do not charge for preprocessing or computations on $G$; only queries to $f$ are counted. We consider both non-adaptive and adaptive testers, with one-sided or two-sided error. A tester is non-adaptive if all queries are fixed in advance (adaptive otherwise), and one-sided if it always accepts monotone functions (two-sided otherwise). Much of the monotonicity-testing literature on both upper and lower bounds focuses on product domains such as the hypercube and hypergrid, where in some settings tight bounds are known [5, 10, 11, 12, 13, 14, 15, 16].
@@
queries, where $\ell$ is the number of comparable pairs, i.e., the number of edges in the transitive closure of $G$. They also show non-adaptive two-sided lower bounds, for Boolean ranges, based on Ruzsa–Szemerédi-type constructions [19], yielding a bound of $n^{\Omega(1/\log\log n)}$. This is still far from the abovementioned upper bound, leaving a large gap.
@@ 1.2 Our results
Our main contribution is improved lower and upper bounds for the query complexity of monotonicity testing on DAGs: we strengthen lower bounds in both non-adaptive and adaptive models, and complement them with improved non-adaptive one-sided upper bounds in sparse regimes. The quantitative statements and their comparison with prior work are summarized in Table 1.
@@
In the adaptive regime, logarithmic lower bounds are known already on the total order (equivalently, path graph) [17]. For constant $\varepsilon$ and infinitely many $n$, we prove an $\Omega(\sqrt{n})$ lower bound on bipartite $n$-vertex DAGs for adaptive one-sided testers (Theorem 4.28).
@@
  • A bound obtained by combining sampling from the transitive reduction and transitive closure: there is a tester with $q=O\!\left(\sqrt{m\,\ell}/(\varepsilon n)\right)$ queries (Theorem 5.1).
@@
These upper bounds can improve the $O(\min\{\sqrt{n/\varepsilon},\ell/(\varepsilon n)\})$ bound in sparse regimes, complementing the lower bounds above. They explicitly exploit both the transitive reduction and the transitive closure, as in work on reachability sparsification and TC-spanner-based testers and reconstruction on posets [6, 7, 28, 31].
@@ 1.3.1 Non-adaptive two-sided lower bound
The classical lower-bound approach for monotonicity testing on general posets, pioneered by Fischer et al. [18], is built around Ruzsa–Szemerédi (RS) graphs: one packs many large induced matchings and uses them to generate hard YES/NO instances. The relevance of matchings is that violated pairs in a matching are vertex-disjoint, so repairing all of them requires changing at least one endpoint per pair, which directly yields a distance-from-monotonicity lower bound. For query lower bounds, matchings also localize evidence: a tester typically needs to query both endpoints of one matched pair to observe informative violations. Inducedness then plays a dual role there. It is what makes the violation pattern “clean” for Boolean-valued functions (extra comparabilities among matching endpoints would immediately create additional violations), and at the same time it becomes the bottleneck: extremal graph theory severely limits how many linear-size induced matchings can coexist in a single graph, which is why RS-based arguments stall far below the $\tilde{\Theta}(\sqrt{n})$ regime.
Our first step is to identify the right analogue of induced matchings for real-valued functions. On bipartite DAGs (all edges oriented left-to-right), the question “can a matching be realized as exactly the set of violated edges?” has a precise answer: this holds if and only if the matching is positive (Section 3.1). Informally, “positive” means there is a vertex weight assignment $w$ such that $w(\ell)+w(r)>0$ exactly on matching edges and $w(\ell)+w(r)\leq 0$ on all other edges. Equivalently, positivity excludes alternating cycles/closed walks on the matched vertices (Theorem 2.3). This motivates the notion of positive-matching RS (PMRS) families, obtained by relaxing RS from induced matchings to positive matchings (Section 3.2). This relaxation is specific to the real-valued setting and bypasses the inducedness barrier in the Fischer et al. argument.
@@
Once a PMRS decomposition is available, the lower bound becomes essentially information-theoretic. Given $s$ edge-disjoint positive matchings $M_{1},\dots,M_{s}$ (each of linear size), we hide a random index $i\in[s]$ and construct two distributions over functions. The YES distribution is monotone, built from a base monotone function with a constant margin on non-matching edges (so small random noise does not create new violations), while the NO distribution is $\varepsilon$-far because it violates a constant fraction of edges in the hidden matching $M_{i}$ (Section 3.3). Crucially, the two distributions have identical single-vertex marginals; they differ only in a local correlation between the two endpoints of a matching edge. Therefore, a non-adaptive algorithm can obtain distinguishing information only if it queries both endpoints of some edge in $M_{i}$. Since the matchings are edge-disjoint, a $q$-query set can fully cover at most $O(q^{2})$ edges, and the probability of hitting the hidden matching index is at most $O(q^{2}/s)$, yielding the $\Omega(\sqrt{s})$ lower bound (Theorem 3.8). A useful technical fact (as with induced matchings) is that positivity is closed under taking submatchings (Section 3.1), which lets us refine each large matching into many smaller positive matchings and obtain the $1/\sqrt{\varepsilon}$ dependence (Sections 3.5 and 3.5).
@@
The remaining challenge is to build explicit graphs with many large edge-disjoint positive matchings. This is precisely where the departure from induced matchings matters: unlike RS graphs, PMRS packings exhibit different extremal behavior and admit simple explicit constructions. Our construction (Section 3.4) uses vertex labels $(x,z)$ with $x\in[N]^{k}$ and an additional coordinate $z\in[N^{2}]$. Each matching is indexed by a shift vector $a$ and consists of translated edges $(x,z)\mapsto(x+a,\,z+\|a\|_{2}^{2})$. The nontrivial point is positivity: we design a weight function $w_{a}$ so that for an edge belonging to shift $b$, the positivity-witness sum $w_{a}(\ell)+w_{a}(r)$ (for edge endpoints $\ell,r$) collapses to a quantity of the form $1/2-\|b-a\|_{2}^{2}$, which is positive exactly when $b=a$ (Theorem 3.9). This quadratic construction separates the shift $a$ from every other shift through the sign of this sum and yields $s=\Omega(n^{k/(k+2)})$ edge-disjoint linear-size positive matchings. Combined with the PMRS $\Rightarrow$ lower-bound argument above, this yields our near-$\sqrt{n/\varepsilon}$ non-adaptive two-sided lower bound.
@@ 1.3.2 Adaptive one-sided lower bound
Violations on a hidden matching and the witness principle.
Moving from non-adaptive to adaptive testers, and from two-sided to one-sided error, requires a different lower-bound argument. In the bipartite-DAG setting (all edges oriented left-to-right), one-sidedness comes with a very strong “witness principle” (Section 4): if the tester has not queried a violated edge, then the partial transcript can always be extended to a globally monotone function. In other words, a one-sided tester is not “estimating distance” indirectly; it is literally searching for a witnessed violation. The challenge is that adaptivity seems to help precisely with such search problems: after observing a few values, the tester can steer future queries toward edges that are more likely to be violated.
Section 4 turns this search problem into an indexing problem. As in the PMRS-based non-adaptive lower bound, we start from a base graph containing many edge-disjoint positive matchings $M_{1},\dots,M_{s}$, and we hide a random index $I\in[s]$. The goal is to generate a distribution over functions such that (i) with high probability the sampled function is $\varepsilon$-far from monotone because it violates almost all edges of the hidden matching $M_{I}$, (ii) edges outside $M_{I}$ are almost never violated (so the tester cannot reject without learning something about $I$), and (iii) even an adaptive algorithm learns only limited information about $I$ per query. The last two requirements are exactly where naive constructions that simply prescribe a violating matching break down: once we condition on an adaptive prefix, the remaining unqueried variables are no longer independent, and a tester might try to create or expose violations on non-matching edges by strategically revealing endpoints.
Our lower-bound argument uses a truncated strongly log-concave prior, namely a Gibbs distribution $\{\mu_{i}\}_{i\in[s]}$ over real-valued labelings $f$ supported on a box. The Hamiltonian has three roles. A quadratic base term
@@
enforces global strong convexity, which later yields stability of conditional marginals under arbitrary transcript conditioning. On top of this, hinge-square potentials implement a monotonicity pattern with margin $\gamma$: for edges $e\notin M_{i}$ we add a penalty $\beta(f(\ell)-f(r)+\gamma)_{+}^{2}$, which favors $f(\ell)-f(r)\leq-\gamma$ (monotone with slack), while for edges $e\in M_{i}$ we add $\beta(\gamma-(f(\ell)-f(r)))_{+}^{2}$, which favors $f(\ell)-f(r)\geq\gamma$ (a violation with margin $\gamma$). Finally, we truncate to a box $B,B]^{V}$ so that the Lipschitz and concentration bounds remain uniform. Under this Gibbs distribution, edges of the hidden matching are violated with overwhelming probability, while edges outside the hidden matching are rarely violated, and this remains true under conditioning on adaptively revealed coordinates.
The rest of the proof has two technical parts: controlling spurious violations and bounding information leakage.
@@
Even if a fresh non-matching edge is very unlikely to be violated unconditionally, an adaptive tester might try to bias it by first querying one endpoint and then deciding whether to query the other. The analysis therefore must be prefix-conditioned: after any transcript prefix, when the tester “closes” a new edge, the conditional probability of witnessing a violation on an edge outside $M_{I}$ must still be tiny. This is handled by a careful one-dimensional reduction on the difference coordinate $d=f(\ell)-f(r)$ for an edge $e=(\ell,r)$: after conditioning on everything except the two endpoints, the induced density of $d$ is log-concave with explicit strong curvature coming from the quadratic term. The hinge term sets a target point ($\gamma/2$ for matching edges and $\gamma/2$ for non-matching edges), and strong convexity translates “the derivative at the target has the right sign” into an exponentially small tail probability of landing on the wrong side.
The subtlety is the external drift contributed by all other incident constraints: under adaptive conditioning, this drift is random and depends on the revealed neighborhood. Section 4 controls this drift at the two critical points $d=\pm\gamma/2$ with a high-probability argument: for $C_{4}$-free base graphs, the relevant neighborhood variables become independent after conditioning on the endpoints, and strong log-concavity gives subgaussian tails for the drift; see Section 4.2. This is paired with a separate bound on the conditional mean of the drift; in the final instantiation on our explicit dense base graph (Section 4.7), this is the main analytic difficulty.
There is a second obstacle specific to the one-sided setting: the box truncation introduces boundary effects, and an adaptive tester might try to push queried values near $\pm B$, where the feasible interval of $d$ shrinks and the clean tail bound could fail. The fix is a deterministic inward bias (Sections 4.1 and 4.1): when $\beta>\lambda$, the hinge potentials create a deterministic force that pushes single-site conditionals away from the dangerous boundary strips. This yields a bound uniform over adaptive strategies on the probability that any query ever lands near $\pm B$ (Section 4.3), and hence bounds the probability that an adaptive strategy can exploit truncation to manufacture spurious witnesses. Putting these together yields a theorem controlling spurious queried violations (Section 4.3): with high probability, any $q$-query adaptive strategy witnesses no violated edge outside the hidden matching unless it hits an exponentially unlikely boundary event.
@@
Even if spurious violations are suppressed, an adaptive tester could hope to learn the hidden index $I$ from the queried values, and then target an edge of $M_{I}$. Section 4 controls this via a mutual-information upper bound $I(I;T)$, where $T$ is the full adaptive transcript. The main analytic ingredient is a “Green-function comparison” inequality for uniformly convex Gibbs measures (Section 4.4): for any prefix conditioning $\tau$, the KL divergence between the one-site conditional distributions under $\mu_{i}(\cdot\mid\tau)$ and $\mu_{j}(\cdot\mid\tau)$ is bounded by a weighted sum of local discrepancies, with weights given by entries of $(\alpha I+2\lambda L_{\tau})^{-1}$, where $L_{\tau}$ is the Laplacian on the still-unqueried vertices after conditioning on the prefix $\tau$. This comparison needs no global weak-dependence or uniqueness assumption: it uses only the strong convexity provided by the quadratic term and survives arbitrary adaptive conditioning. A simple row-sum bound on this Green function then gives a per-query leakage bound $\kappa_{Q}$, leading to the budgeted bound $I(I;T)\leq q\kappa_{Q}$ uniformly over all adaptive strategies with $q\leq Q$ (Section 4.4).
With control of spurious queried violations and leakage in place, the lower bound becomes information-theoretic and well suited to one-sided testing. Because the matchings are edge-disjoint, any witnessed violated edge in $\bigcup_{i}M_{i}$ uniquely identifies the hidden index $I$. By the one-sided witness principle, rejection implies the transcript contains a witnessed violated edge; by Section 4.3, with high probability this violated edge lies in $M_{I}$. Thus any successful tester yields a decoder that recovers $I$ from $T$ with constant probability, and a standard Fano argument (Section 4.5) forces $I(I;T)\gtrsim\log s$. Comparing $\log s$ to the leakage upper bound $q\kappa_{Q}$ yields a lower bound for testers with query budget $q\leq Q$ (Theorem 4.22): below a certain budget threshold, no adaptive one-sided tester can simultaneously (a) keep spurious witnesses rare and (b) accumulate enough information to find a true matching-edge violation reliably.
@@
Section 4.7 applies this argument to an explicit dense base graph: a $C_{4}$-free, almost-regular subgraph of the shift-matching construction (Section 4.7) with $s=\Theta(\sqrt{n})$ disjoint positive matchings. In this regime the auxiliary probabilities (matching-edge non-violation, spurious violations, boundary failures, and mean-drift failures) can all be driven superpolynomially small by setting the margin $\gamma$ to grow like $\sqrt{\log Q}$, while the per-query leakage remains on the order of $(\log Q)/\sqrt{n}$. Choosing $Q\asymp\sqrt{n}$ then yields the claimed $\Omega(\sqrt{n})$ adaptive one-sided lower bound for a constant $\varepsilon$ (Theorem 4.28).